### v1

@@ 4.3. Multi-task Learning for Recommendation
In multi-task learning we aim to learn multiple tasks simultaneously and attempt to maximize performance on a subset or all of them, sometimes striking balance or trade-off between tasks. This is practically interesting to a marketplace as users can perform many different actions to any item, therefore correctly anticipating the chance of each action can invite additional user experience optimizations such as showing appealing alternative items, providing economical incentives, etc.
@@
For example benchmark we chose two competitive models, the MMOE (Ma et al., 2018b) based on mixture of experts (Jacobs et al., 1991) with additional gating mechanisms for all tasks, as well as ESMM (Ma et al., 2018a) which models multiple tasks with explicit inter-task dependent relationships and a factorized loss function. In addition to the default two-task configuration, we also evaluated MMOE on single task (item_view-only and item_like-only) to show the difference against having two tasks simultaneously trained. Other approaches such as ESMC (Jiang et al., 2023), RMTL (Jiang et al., 2023), PEPNet (Chang et al., 2023) were omitted due to time and resource constraints. Similar to the CTR prediction (Section 4.1) setup, here we also omit extensive hyperparameter search for time and resource reasons. See Table 10 for hyperparameter configuration details. All executions in this subsection were ran on an Nvidia L4 GPU equipped linux system with 16 cores, 128GB RAM on Google Cloud Platform. See Table 6 for the results using area under the curve (AUC) as the metric.
@@
We observe that MMOE performs marginally better when using both tasks simultaneously than with either task in isolation, at the expense of overall lengthened training time. ESMM outperformed the rest by insignificant amount on MerRec, whereas more pronounced performance gaps were observed on other benchmark datasets (e.g. (Yuan et al., 2022)). This shows that MerRec has a highly diverse and challenging environment to tune and explore these types of models.
@@
Model view AUC like AUC Train+Val Time (Hrs) VRAM (GB)
Only view 0.709 N/A 16.3 3.93
Only like N/A 0.709 16.6 3.93
MMOE 0.712 0.713 22.61 3.94
ESMM 0.713 0.715 23.7 3.94
@@ 5. Conclusion
In conclusion, sourced from the Mercari platform, we propose the large-scale [Uncaptioned image] dataset for Consumer-to-Consumer (C2C) recommendation. Our exploration into the MerRec dataset has highlighted the evolving role of recommendation systems in e-commerce, especially within the marketplace. The introduction of the Mercatran model, tailored to address the unique challenges of C2C environments, represents a significant advancement in recommendation system research. Through tasks like click-through rate prediction, session-based recommendation and multi-task learning for user action prediction, we’ve showcased the potential of the MerRec dataset and Mercatran to enhance user experience by accommodating the dynamic nature of C2C interactions.

### v2

@@ 4.3. Multi-task Learning for Recommendation
In multi-task learning we aim to learn multiple tasks simultaneously and attempt to maximize performance on a subset or all of them, sometimes striking balance or trade-off between tasks. This is practically interesting to a marketplace as users can perform many different actions to any item, therefore correctly anticipating the chance of each action can invite additional downstream user experience optimizations such as showing appealing alternative items, providing economical incentives, suggesting user to negotiate a counter offer or saving the item for consideration later depending on the business needs.
@@
For example benchmark we chose two competitive classical models, for example OMoE (Ma et al., 2018b) and MMOE (Ma et al., 2018b) based on mixture of experts (Jacobs et al., 1991) with additional gating mechanisms for all tasks, ESMM (Ma et al., 2018a) which models multiple tasks with explicit inter-task dependent relationships and a factorized loss function, as well as current state-of-the-art approaches such as AITM (Xi et al., 2021), PLE (Tang et al., 2020) and STEM (Su et al., 2024). In addition to the default two-task configuration, we also evaluated MMOE on single task (item_view-only and item_like-only) to show the difference against having two tasks simultaneously trained. Other methods such as ESMC (Jiang et al., 2023), RMTL (Jiang et al., 2023), PEPNet (Chang et al., 2023) were omitted due to time and resource constraints. The ME-MMOE and ME-PLE competitive baselines introduced by the STEM (Su et al., 2024) paper were omitted here as certain important implementation detail were not publicly available.
Similar to the CTR prediction (Section 4.1) setup, here we also omit extensive hyperparameter search for time and resource reasons. See Table 12 and Appendix Section A.1 for hyperparameter tuning and configuration details. All executions in this subsection were ran on an Nvidia L4 GPU equipped linux system with 16 cores, 128GB RAM on Google Cloud Platform. See Table 7 for the results using area under the curve (AUC) as the metric.
@@
We observe that MMOE performs marginally better when using both tasks simultaneously than with either task in isolation, at the expense of overall lengthened training time. ESMM outperformed MMOE by insignificant amount on MerRec, whereas more pronounced performance gaps were observed on other benchmark datasets (e.g. (Yuan et al., 2022)). AITM, PLE, OMoE and STEM generally performed better and were easier to trigger early stopping during training, and there is a trend of performing better on the view task relative to the like task, whereas MMOE and ESMM performed better on the like task relative to the view task. STEM performed marginally worse than AITM, PLE and OMoE despite being the SoTA reported on other datasets (Su et al., 2024). This shows that MerRec has a highly diverse and challenging environment to tune and explore these types of models.
@@
Model view AUC like AUC Log Loss (View + Like) Train+Val Time (Hrs) VRAM (GB)
Only view 0.709 N/A 0.395 16.3 3.93
Only like N/A 0.709 0.354 16.6 3.93
MMOE 0.712 0.713 0.744 22.61 3.94
ESMM 0.713 0.715 0.744 23.7 3.94
AITM 0.773 0.736 0.684 2.85 3.26
PLE 0.773 0.736 0.686 4.38 3.39
OMoE 0.773 0.737 0.684 3.3 4.17
STEM 0.772 0.736 0.683 9.68 11.27
4.4. Inference Acceleration for Recommendation
4.4.1. Task Description.
In real world use cases, recommendation predictions on marketplace can be pre-generated offline or live-generated online in real time. Regardless of the setup, we would like to study whether sequential recommendation models can enjoy inference speed ups to improve system efficiency [^11] Experiments implemented here in a way that is comparable to past relevant academic recommendation literature, but inefficient compared to online production system standards. There are significant rooms for improvement in the program, and this preliminary study is only intended to examine potential relative speedup effects rather than serving as a basis to form generalized claims..
4.4.2. Dataset Setup.
In this section we reuse exactly the same dataset format as Section 4.2, down to the same seed, split ratio, etc.
4.4.3. Baseline Setup.
Some work have shown that certain recommendation models can enjoy inference speedups while maintaining a similarly accurate performance. Chen et al. (Chen et al., 2021) proposed the SkipRec framework which attaches a personalized policy network to learn a user-specific set of binary skip-or-keep decisions over the residual blocks of a modified NextItNet model backbone. They argue that by learning a user-specific block-selection policy, the overall computation is adapted to how “easy” or “hard” a user action sequence needs to be computed. SkipRec was introduced with two policy network design options based on the Gumbel-Softmax trick over all such binary decisions, or a self-critical sequence training (SCST) (Rennie et al., 2017) based reinforcement learning approach with curriculum learning (Bengio, 2013; Wu et al., 2018) that incentivizes using minimal active blocks. Following the demonstration by SkipRec (Chen et al., 2021) and TenRec (Yuan et al., 2022), we study its effect on NextItRec (Yuan et al., 2019) and SASRec (Kang and McAuley, 2018) using the Gumbel-softmax variant of SkipRec. All hardware configuration along with hyperparameters are identical to those used for Section 4.2 (except prediction batch size is now changed to 1 instead of 32, a practice we followed from (Yuan et al., 2022)) and we show our experimental results at Table 8. Please note that the total inference test wall time calculated is affected by numerous factors, such as batch size, test set size, etc.
Table 8. Results of Inference Acceleration (IAR).
Model nDCG @5 nDCG @20 Recall @5 Recall @20 Test Inference Time (Hrs) VRAM (GB)
NextItNet 0.240 0.279 0.326 0.461 22.45 16.89
Skip-NextItNet 0.195 0.229 0.265 0.383 22.45 15.11
SASRec 0.152 0.183 0.207 0.314 21.84 12.51
Skip-SASRec 0.209 0.242 0.281 0.397 21.64 14.40
4.4.4. Implications.
We observed that when experimenting on the MerRec dataset, the overall speedup was small or practically negligible, but SkipRec’s impact on the backbone’s performance can be inconsistent: On NextItNet it can be too aggressive in skipping blocks which caused a degradation in NDCG and recall, while on SASRec it shortened the prediction time while boosted the NDCG and recall by significant amounts, indicating that the benefit or trade-offs of attaching SkipRec onto a sequential recommendation model backbone may be case dependent under the highly diverse MerRec dataset.
@@ 5. Conclusion
In conclusion, sourced from the Mercari platform, we propose the large-scale [Uncaptioned image] dataset for Consumer-to-Consumer (C2C) recommendation. Our exploration into the MerRec dataset has highlighted the evolving role of recommendation systems in e-commerce, especially within the marketplace. The introduction of the Mercatran model, tailored to address the unique challenges of C2C environments, represents a significant advancement in recommendation system research. Through tasks like click-through rate prediction, session-based recommendation, multi-task learning for user action prediction and inference acceleration in recommendation models, we’ve showcased the potential of the MerRec dataset and Mercatran to enhance user experience by accommodating the dynamic nature of C2C interactions.
@@ Appendix A Experiment Configuration
A.1. CTR, MTL, SBR and IAR Hyperparameter Tuning (Excluding Mercatran)
In CTR, MTL, SBR and IAR experiments (excluding Mercatran), we use a narrower set of candidate values per hyperparameter than those reported by the cited references, by picking the commonly reported best values (e.g. (Yuan et al., 2022; Su et al., 2024; Ma et al., 2018b) as well as most of the individual model-specific papers) and run grid search on them. For example, in the Multi-task learning section we observed from (Yuan et al., 2022; Su et al., 2024; Ma et al., 2018b) that most of the models often have either 16 or 32 for their embedding dimension, so these two values would become part of the grid search. We then repeatedly collect values like these for hidden dimension (e.g. for MLP layers in certain models like MMOE (Ma et al., 2018b) and OMoE (Ma et al., 2018b)), and other factors such as learning rate (e.g. 1e-3, 5e-4, 1e-4) for the Adam (Kingma and Ba, 2015) optimizer and batch size (e.g. 1024, 2048, 4096). Other model-specific hyperparameters (such as the attention factor for AFM (Xiao et al., 2017)) are set to their single suggested value to limit the total amount of computation used.
One detail to note, is that we treat the embedding dimension slightly special, because the standard practice we observed is to use the same embedding dimension for all models on the same experiment for fairness, hence we try to manually look at the metrics across the spectrum on which embedding dimension brings higher values on average, then choose the better one to base the analysis on. One of the observations we had was that the embedding size tends to have a bigger impact on metric performance and training efficiency than most of the other hyperparameters, and selecting the other hyperparameters to be mostly the same values across models did not cause any significant discrepancies or unbalancing to the evaluation, as the performance variations across different other configurations were usually ¡1% apart once seed is fixed. Similar phenomenons were also observed in other CTR (Yuan et al., 2022), MTL (Yuan et al., 2022; Su et al., 2024; Ma et al., 2018b), and SBR (Yuan et al., 2022) studies to an extent.
A.2. Mercatran Hyperparameter Tuning
For Mercatran the hyperparameters are selected based on a variety of factors such as recall and nDCG as reported in the main manuscript but also cost. Given that the model is used to embed user interactions as well as to-be-indexed items coming into the platform, the smaller the embedding size, the smaller the storage needed. Hence we performed a grid search over the embedding size (with the goal to reduce the size) while also measuring the impact on the metrics simultaneously. For embedding size, we started with 512 and ultimately stopped at 64 below which we experienced significant degradation in our primary metrics. We used the same strategy for other parameters that also affect the cost such as number of stacked transformer blocks, sequence length, etc.
For parameters not related to the cost such as learning rate, batch size etc. as mentioned above, we perform grid search on most commonly reported parameters.